OpenStudy (anonymous):

How do i go about solving this question 10?

OpenStudy (anonymous):

http://www.math.uh.edu/~pamb/1431EMCF%2012.pdf

OpenStudy (anonymous):

what is the derivative of f(x) = sin^2(x) ?

OpenStudy (anonymous):

you should have a formula for that derivative somewhere :)

OpenStudy (anonymous):

it is 2sinx cos x

OpenStudy (anonymous):

I dont know where to go from that besides putting in the pi over 4 and using the unit circle points..

OpenStudy (anonymous):

oh, well you're almost there... If you have that as the derivative... that expression is f'(x) f ' (x) = 2(sin (x) )(cos (x) ) and you need to evaluate at x = pi / 4

OpenStudy (anonymous):

so f ' ( pi/4) = 2 sin (pi/4) cos (pi / 4)

OpenStudy (anonymous):

2(root2/2)(root2/2)

OpenStudy (anonymous):

then i dont understand what im supposed to do next

OpenStudy (anonymous):

\[2\frac{ \sqrt{2} }{ 2 }\frac{ \sqrt{2} }{ 2 } = 2 \frac{ \sqrt{2}^{2} }{ 4 } = 2\frac{ (2) }{ 4 }\]

OpenStudy (anonymous):

**mind blown*** why ? my teacher gave no examples like this btw. so im a noob

OpenStudy (anonymous):

:) I love "mind status: blown" it's a sign of a curious (= healthy) person !! But no kidding... wish they'd teach you instead of leaving so much for you to discover on your own!!

OpenStudy (anonymous):

oh wait i see it. i should get some sleep lol.

OpenStudy (anonymous):

ahaha im silly i thought the = sign was a plus sign so i was like O.e sorcery! lol buuut i see it now. so it would just be 1

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

yeah :) we both need some sleep :)

OpenStudy (anonymous):

pretty sure that "sorcery" explains it all... :) that will laugh me off to sleep :)